Jim has 19 quarters.Mathew has 18 pennies.Jayleen has 10 nickels.Daznet has 18 dimes. Who much coins are there in all ?
Answer by Screamgirl(4)

19 quarters,18 pennies, 10 nickels, 18 dimes
19+18=37+10=47+18=65
Answer: 65 coin together
The amount they have is $7.23